1. Two conducting plates with a length of 1 m and a width of 0.5 m has polyethylene between them. What is the capacitance if the capacitor is 1 m high?



1
Expert's answer
2022-03-16T18:28:17-0400

The capacitance of the parallel plates capacitor

"C=\frac{\kappa\epsilon_0 A}{d}"

"C=\frac{2.25*8.85*10^{-12}*1}{0.5}=4.0*10^{-11}\:\rm F"